3) The reaction NO2 (g) NO (g) N20 0g)+ O2 (g) reached equilibrium at a certain high temperature. Originally, the reaction vessel contained the following initial concentrations: 0.184 M N20, 0.377 M O2, 0.0560 M NO2 and 0.294 M NO. The concentration of NO2, the only colored gas in the mixture, was monitored by following the intensity of the color. At equilibrium, the NO2 concentration had become 0.118 M. What is the value of Kc for the reaction at this temperature? Also determine the equilibrium concentrations of N20, o2, and NO.
Answer #1

[NO2]

[NO2] [NO]

[N2O]

[O2]

I

I 0.0560

0.294

0.184

0.377

C

+x

+x

-x

-x

E

0.0560 + x

0.294 + x

0.184 – x

0.377 – x

[NO2] = 0.118 M = 0.0560 + x at equilibrium.

Solving

x = 0.062 M.

Using the ICE table ------------

that the equilibrium concentrations are;

[NO] = 0.356 M,

[N2O] = 0.122 M

And

[O2] = 0.315 M.

Kc = [N2O] [O2] /[N2O] [NO] = (0.122)(0.315) /(0.118)(0.356) =0.915-----ANSWER